CSAT3 redeployment

As we continue to spin up gas flux measurements, CSAT3s will be installed as needed.

Currently, CSAT3 S/N 2678 has been reinstalled in the LEO east bay as of 12/12/2023.

EX81 flow meters zero reset

All three ex81 zeros were reset. This was done between 1-2pm.

We have been seeing non-zero flow rates when there is no flow.

Currently, Center bay seems to be metering less flow with quick initial test with MP2000.

Further evaluation will be done.

Biweekly rain modification due to power fluctuation

On 9/12/2023, the non-experimental biweekly rain was interrupted by multiple power disruptions by the utility power company. We chose to discontinue the rain and resume the remaining rain duration on 9/13/2023.

On 9/12/2023, irrigation of multi one and multi two rain circuits of approximately 3 hours 15 minutes (0830-11.45) was completed. The remaining 3 hours 45 minutes for a net duration of 7 hours was completed on 9/13/2023 0900-1245.

NovaLynx calibration

The LEO NovaLynx 260-2501-A 8'' Rain Gauges were calibrated on September 26-28, 2022.

For details on the calibration please see the novalynx_calibration_2022_2.xlsx.

Calibration values were updated on 11:53am 10/3/2022.

 

Testing wiki email

This is a test of sending emails

Test new blog

This is just a test for GLOG content visualization.

Biosphere 2 logo

 

LEO center SDI12 subsoil sensor offline

LEO Center SDI12 subsoil sensors when offline after 3/1/2022 at 9pm.

The power supply to these sensors was found to be dead.

It was replaced on 3/14/2022 around 12:40pm.

Since there was no power to the sensors, data at this offline period will not exist.

Higher Infiltration Rate

On February 1, 2022 the standard bi-weekly rain was altered to explore the impact of the developed surface crust at higher infiltration rates.

The standard schedule is: rain from 7:45am to 2:45pm with Multi2 at approximately 13mm/h.

The plan was to cut the rain one hour short and rain the previous hour at double intensity (7:45am-12:45pm with Multi2 at ~13mm/h and 12:45pm-13:45pm with Multi1 and Multi2 at ~26mm/h).

Overland flow started shortly after switching to the higher infiltration rate. Rain was stopped at ~13:03pm. 

As of right now, the rain rate should not exceed 13mm/h.

Short LEO rain

On 3/8/2021, the standard bi-weekly rain from 7:45am to 2:45pm was cut short at 1pm.

This was done to prevent our VFD pumps from running dry as there was insufficient water in the main LEO irrigation tanks.

The auxiliary tanks water was not pumped into the main tanks early enough.

LEO mast kept raised short term

To facilitate various work on the hillslope and due to lack of personnel, the LEO masts have been kept raised since 6/8/2021. This is a short term measure.

Tags:
Created by Michael Sibayan on 2022/05/03 16:48
  
This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 13.10.5 - Documentation